You wish to prepare 100 mL of a buffer solution that is 0.025 M in carbonic acid (H2CO3). The pKa of H2CO3 is 6.351. Calculate the amount of H2CO3, in grams, that you will need to weigh out.
Answer #1

Since you haven't mentioned pH of buffer solution

We are going to calculate it by H2CO3 concentration.

pH of 0.025 M H2CO3

pKa = -log[Ka] = 6.351

Ka = 4.46 x 10^-7

H2CO3 <==> H+ + HCO3-

let x be the amount of dissociation

Ka = [H+][HCO3-]/[H2CO3]

4.46 x 10^-7 = x^2/0.025

x = [H+] = 1.056 x 10^-4 M

pH = -log[H+] = 3.98 [Substituting this value with a differnt number would change all the further calculation results]

Using Hendersen-Hasselbalck equation,

pH = pKa + log)[base]/[acid])

3.98 = 6.351 + log([HCO3-]/[0.025])

[HCO3-]/[H2CO3] = 0.0042

[H2CO3] = (1/1.0042) x (0.025/0.1) = 0.025 M

[HCO3-] = (0.0042/1.0042) x (0.025/0.1) = 1.046 x 10^-4 M

grams of H2CO3 required = 0.025 x 0.1 x 62.03 = 0.155 g

pl. note. If you have a pH value for buffer, use Hendersen-Hasselbalck equation directly and follow the procedure with your calculated value accordingly.
Answer #2

Hi, I think that may be the question is incomplete, because it doesn